OpenStudy (anonymous):

why is the derivative of y=sqrt(x), y'=1/2x^-1/2?

OpenStudy (anonymous):

are you asking how to get this from scratch? \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\]?

OpenStudy (anonymous):

if you know the power rule it is a straightforward application

OpenStudy (anonymous):

yea that is what im asking from scratch like why ise power the same still?

OpenStudy (anonymous):

why is the power the same still?*

OpenStudy (anonymous):

sqrt(x) = x^1/2

OpenStudy (anonymous):

well power rule says \[\frac{d}{dx}x^r=rx^{r-1}\]

OpenStudy (anonymous):

1 * 1/2 = 1/2

OpenStudy (anonymous):

and \[\sqrt{x}=x^{\frac{1}{2}}\]

OpenStudy (anonymous):

1x^1/2

OpenStudy (anonymous):

so power rule give it to you right away

OpenStudy (anonymous):

\[\color{red}{\text{right myininaya?}}\]

OpenStudy (anonymous):

power rules of x^n = nx^(n-1)

OpenStudy (anonymous):

yea i have to subtract by one..i did it on the calculator and got the answer

OpenStudy (anonymous):

but you can always compute \[\lim_{\rightarrow 0}\frac{\sqrt{x+h}-\sqrt{x}}{h}\] and get the same thing

OpenStudy (anonymous):

yeah, but it simplifies to d/dx x^n = nx^(n-1)

OpenStudy (anonymous):

btw this function is very common, so it is best just to remember it and not use the power rule each time. kind of like memorizing what 7 times 8 is

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{\sqrt{x+h}-\sqrt{x}}{h}=\lim_{h \rightarrow 0}\frac{\sqrt{x+h}-\sqrt{x}}{h}*\frac{\sqrt{x+h}+\sqrt{x}}{\sqrt{x+h}+\sqrt{x}}\] \[=\lim_{h \rightarrow 0}\frac{x+h-x}{h(\sqrt{x+h}+\sqrt{x}}=\lim_{h \rightarrow 0}\frac{1}{\sqrt{x+h}+\sqrt{x}}=\frac{1}{\sqrt{x+0}+\sqrt{x}}\] \[=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

yes, but proving it to yourself first makes you feel confident about it

OpenStudy (anonymous):

what myininaya said. but also if you are taking calc, you should just remember that the derivative of the square root of x is one over 2 root x. it will help you for sure

myininaya (myininaya):

oops i forgot to close one of my parenthesis :(

OpenStudy (anonymous):

ohhh okay bt when i need to show my work the power rule will be in memory

OpenStudy (anonymous):

also, remember x^(-1/2) = 1/(x^(1/2))